The price of a new car is $40,000. Assume that an individual makes a down payment of 25% toward the purchase of the car and secures financing for the balance at the rate of 9%/year compounded monthly. (Round your answers to the nearest cent.)

(a) What monthly payment will she be required to make if the car is financed over a period of 24 months? Over a period of 72 months?

24 months     $
72 months     $


(b) What will the interest charges be if she elects the 24-month plan? The 72-month plan?

24-month plan     $
72-month plan     $
Answer #1

A B 1 a) 24 months 2 72 months $1,370.54 =PMT(9%/12,24,-40000*(1-25%)) $ 540.77 =PMT(9%/12,72,-40000*(1-25%) | 4 b) |24 month

*Please rate thumbs up